What is the smallest whole number that has a remainder of 1 when divided by 4, a remainder of 1 when divided by 3, and a remainder of 2 when divided by 5?
Answer by RAY100(1637)

5,,,,4,,,,7,,,,
.
.
check,,,5/4=1R1,,,,4/3=1R1,,,,7/5=1R2
.
.